# Local Mixtures of Experts: Essentially Free Test-Time Training via Model Merging

### v1

@@ Abstract
Mixture of expert (MoE) models are a promising approach to increasing model capacity without increasing inference cost, and are core components of many state-of-the-art language models. However, current MoE models typically use only few experts due to prohibitive training and inference cost. We propose Test-Time Model Merging (TTMM) which scales the MoE paradigm to an order of magnitude more experts and uses model merging to avoid almost any test-time overhead. We show that TTMM is an approximation of test-time training (TTT), which fine-tunes an expert model for each prediction task, i.e., prompt. TTT has recently been shown to significantly improve language models, but is computationally expensive. We find that performance of TTMM improves with more experts and approaches the performance of TTT. Moreover, we find that with a 1B parameter base model, TTMM is more than 100×100$\times$100 × faster than TTT at test-time by amortizing the cost of TTT at train-time. Thus, TTMM offers a promising cost-effective approach to scale test-time training. [^1] The implementation is available at: https://github.com/rbertolissi/ttmerge
Figure 1: Accuracy gains of TTMM when scaling test-time compute (left) and pre-training compute (right). The normalized perplexity between base model and TTT model is averaged across evaluation datasets and models. Left: We measure language modeling ability against the time to generate 100 tokens. TTMM approaches the performance of TTT without almost any test-time overhead when increasing the number of active experts. Ensembling rather than merging experts improves performance, but at a much higher test-time cost. Standard fine-tuning on the training data improves performance without any test-time overhead, but does not yield close to the performance of TTT. Right: We measure the performance of TTMM against the total number of experts. Whereas common MoE models use few experts, TTMM scales to an order of magnitude more experts, improving performance with more experts. Scaling the number of experts in TTMM does not affect training FLOPs or required GPU memory. [^3] Scaling the number of experts increases CPU host memory which is cheaper than GPU memory. Training the individual experts is embarrassingly parallelizable.
@@ 1 Introduction
The standard paradigm of machine learning separates training and testing. In training, we learn a model by inductively extracting general rules from data, while in testing, we apply this model to new unseen data. We study an alternative transductive paradigm called test-time training (TTT, Sun et al., 2020) where we use a specific expert model for each prediction task. Variations of this paradigm have been studied since the inception of the machine learning field (Cleveland, 1979; Cleveland & Devlin, 1988; Atkeson et al., 1997; Bottou & Vapnik, 1992). More recently, fine-tuning large pre-trained neural networks at test-time has regained interest in computer vision (e.g., Jain & Learned-Miller, 2011; Shocher et al., 2018; Sun et al., 2020; Gandelsman et al., 2021; Ruiz et al., 2023) and language modeling (Krause et al., 2018; 2019; Hardt & Sun, 2024; Sun et al., 2024; Hübotter et al., 2025). Hardt & Sun (2024) and Hübotter et al. (2025) fine-tune a pre-trained language model on data related to the prompt, and show that this can significantly improve language modeling performance. Beyond pure language modeling, TTT has been instrumental in state-of-the-art approaches to abstract reasoning (Akyürek et al., 2024), video generation (Dalal et al., 2025), and also been shown to improve reasoning language models (Zuo et al., 2025; Simonds & Yoshiyama, 2025).
@@
In our experiments, TTMM approaches the language modeling performance of TTT without incurring almost any significant compute or memory cost (cf. Footnote 3). Figure 2 illustrates the difference between TTT and TTMM.
@@
We contribute the following key findings, summarized in Footnote 3:
@@
    TTMM outperforms multitask model merging: We find that training and merging many expert models — even within a single domain such as “coding in Python” — outperforms training and merging few models specialized on broader tasks.
@@ 2 Related Work
Combining the predictions of multiple models has been a long-standing practice in machine learning, e.g., in the form of majority voting (Wang et al., 2023) or by ensembling of predictions (Dietterich, 2000; Wortsman et al., 2022). However, merging predictions of many neural networks is computationally inefficient since it requires a separate forward pass for each model. Instead, TTMM merges in parameter-space, and only requires a single forward pass of the merged model, resulting in a negligible increase of compute and memory cost compared to inferencing a single model. The idea of merging multiple models by averaging their parameters (not their predictions!) is grounded in the surprising phenomenon of mode-connectivity (Garipov et al., 2018), which observes that modes in ensembles of models are connected by paths of small loss. This has motivated a variety of model merging techniques, including stochastic weight averaging (Izmailov et al., 2018) and model soups (Wortsman et al., 2022). Recently, model merging has been studied predominantly in the context of merging few expert models with the aim of retaining performance on their respective meta-tasks (Ilharco et al., 2023; Matena & Raffel, 2022; Jin et al., 2023; Ainsworth et al., 2023). Merging model parameters may lead to interference, which has been a key focus of recent work (Yadav et al., 2023; Yang et al., 2024c; Tang et al., 2024b; Daheim et al., 2024; Yu et al., 2024; Jung et al., 2024).
@@
While the above literature focuses on multitask learning with typically only a handful of expert models, TTMM clusters the training data into many local neighborhoods and trains a separate local model for each cluster. Like TTMM, Branch-Train-Merge (BTM, Li et al., 2022; Gururangan et al., 2023) clusters the training data into local neighborhoods, trains a separate local model for each cluster, and dynamically selects a subset of these models at test-time. However, unlike BTM, TTMM uses an order of magnitude more local models and merges their parameters instead of their predictions, leading to substantially faster inference. Li et al. (2024) have recently shown that training and dynamically selecting from millions of LoRA adapters can further improve performance and factual accuracy of LLMs, yet unlike TTMM, they also do not use parameter merging.
@@
The mixture of experts (MoE, Shazeer et al., 2017) paradigm increases model capacity without increasing inference cost by introducing conditional routing of inputs to subsets of learnable parameters (i.e., “experts”). Training MoE models is challenging due to the need for joint training of the experts and the router on substantial multitask data. Moreover, MoE models evaluate all active experts individually which can still require substantial memory. This in turn implies that the number of experts is typically limited to a small number (commonly at most $10$ experts). Recent studies have focused on challenges such as reducing synchronization (Sukhbaatar et al., 2024; Zhang et al., 2024; 2025), load balancing of experts (Clark et al., 2022; Zhou et al., 2022), and improving expert specialization (Dai et al., 2024). In contrast, TTMM approximates the MoE paradigm by embarrassingly parallelizing the training of the local models and merging model parameters at test-time, which permits scaling to many orders of magnitude more experts and which enables approximation of the “maximally local” TTT.
@@ 3 TTMM: Test-Time Model Merging
Our proposed method TTMM has two main stages: train-time and test-time. At train-time, TTMM clusters the training data into many local neighborhoods and trains a separate expert model for each cluster. Then at test-time, TTMM dynamically composes these expert models to form a single task-specific model. We discuss both stages in more detail below.
@@
3.1 Test-Time Efficiency
In contrast to MoE models, the number of experts in TTMM is typically too large for all experts to fit in GPU memory at once. For example, with Llama-3.2-1B as base model, each LoRA adapter has size approximately $172$ MB, and with $1000$ clusters, the expert models would require prohibitive $172$ GB of GPU memory. Instead, we dynamically load active experts into GPU memory and merge their parameters before inference.
Figure 4: TTMM has negligible compute overhead at test-time. Left: Inference cost of ensembling and TTMM with $10$ (dark) and $3$ (light) active experts, respectively. TTMM with $10$ experts costs around $20$ generated tokens, while with $3$ experts, TTMM costs only about $10$ generated tokens. In contrast, ensembling has multiplicative overhead that grows with the number of active experts. Middle: Test-time overhead of TTMM compared to inference with the base model. When using up to $10$ experts, the test-time overhead is small. Right: Number of active experts with a given temperature $\beta$ and sparsity $\tau$ on the Wikipedia dataset with $K=1000$ experts. Setting: We use Llama-3.2-1B as base model and consider LoRA adapters with rank $64$. Each LoRA adapter has size $\approx\!172$ MB. For loading, we assume a CPU to GPU bandwidth of $50$ GB/s. For computing embeddings, we use the all-mpnet-base-v2 model. We measure selection and merging on an RTX 4090 GPU with $10$ seeds.
The test-time overhead of TTMM consists therefore of the time required to compute the merging coefficients (select), the time to load active experts from CPU to GPU memory (load), and the time to merge the expert parameters (merge). In Figure 4 (middle), we benchmark this overhead across varying numbers of active experts. On our infrastructure and with $10$ active experts, selection takes roughly $5$ milliseconds, loading takes around $30$ milliseconds, and merging takes approximately $80$ milliseconds. We use PyTorch’s einsum (Paszke, 2019) for computing the merged LoRA adapter $\Delta\boldsymbol{W}=\sum_{k}w_{k}\boldsymbol{B}_{k}\boldsymbol{A}_{k}$ (cf. Figure 7 in the appendix), with $\boldsymbol{A}_{k}$ and $\boldsymbol{B}_{k}$ the low-rank factors of the LoRA adapter $\boldsymbol{\theta}_{k}$. This is around $20\%$ faster than the naive implementation with a for-loop over the experts, without any additional memory overhead. We anticipate that test-time overhead may be reduced further by interleaving loading and merging by overlapping computation and communication (see, e.g., Tang et al., 2025).
@@ 3.1 Test-Time Efficiency
In Figure 4 (middle), we evaluate the practical test-time overhead of TTMM for language generation. TTMM incurs only a small fixed overhead which, when merging $10$ experts, is equal to generating around $20$ tokens. When generating more than $100$ tokens, the overhead of TTMM compared to inference with the base model is negligible. In contrast, ensembling has a multiplicative overhead that grows with the number of active experts. Comparing to TTT, a single test-time gradient step with Llama-3.2-1B takes approximately $150$ milliseconds (Hübotter et al., 2025). With $100$ test-time gradient steps, the total test-time overhead (without any data selection and loading) is already around $15$ seconds. In contrast, TTMM with $10$ experts costs around $115$ milliseconds which is a more than 125×125$\times$125 × speedup compared to TTT.
@@ 3.2 Test-Time Model Merging approximates Test-Time Training
Suppose the neural network $f(\boldsymbol{x};\boldsymbol{\theta})$ is $L$-Lipschitz in $\boldsymbol{\theta}$ and that the loss is $G$-smooth in $\boldsymbol{x}$, with any $L,G>0$. Further, suppose we adapt models via single-step gradient descent with step size $\eta>0$ from a shared initialization $\boldsymbol{\theta}$. For any prompt $\boldsymbol{x^{\star}}\in\mathcal{X}$ and $N\geq 1$, let $\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}$ be the TTT model trained on the set $\mathcal{D}_{\boldsymbol{x^{\star}}}$ of the $N$ nearest neighbors to $\boldsymbol{\phi}(\boldsymbol{x^{\star}})$ in $\mathcal{D}$. Let $\boldsymbol{\theta}^{\prime}$ be the expert model trained on any $\mathcal{D}^{\prime}\subseteq\mathcal{D}$. Then, if $\mathcal{D}^{\prime}$ contains at least one nearest neighbor from $\mathcal{D}_{\boldsymbol{x^{\star}}}$,
To understand the intuition behind this proposition, consider an expert model trained on a cluster $\mathcal{D}^{\prime}$ that contains parts of the local neighborhood of the prompt embedding (for this, the entire clustering must cover the dataset). Then, the above proposition shows that this expert model is a good approximation of the TTT model if the diameter of $\mathcal{D}^{\prime}$ is also small. Consequently, this motivates the approximation of TTT by TTMM with many expert models covering the entire dataset, each trained on a small cluster. Training expert models on this partitioning of the dataset does not cost more FLOPs or GPU memory than training a single model on the entire dataset. This is in contrast to TTT, which considers a separate expert model for each possible recombination of the training data into subsets of size $N$.
@@
Instead of selecting the expert model by minimizing the sum of distances of individual data points to the prompt embedding (i.e., $\smash{\sum_{\boldsymbol{x}\in\mathcal{D}_{k}}\left\|\boldsymbol{\phi}(\boldsymbol{x^{\star}})-\boldsymbol{\phi}(\boldsymbol{x})\right\|}$), TTMM summarizes each expert model by its centroid $\smash{\nicefrac{{1}}{{|\mathcal{D}_{k}|}}\sum_{\boldsymbol{x}\in\mathcal{D}_{k}}\boldsymbol{\phi}(\boldsymbol{x})}$. Selecting the expert with the closest centroid to the prompt embedding further speeds up inference, since we only need to compute $K$ distances as opposed to the distances to all training data embeddings. This simplification does not come for free: Consider the one-dimensional set of embeddings $\{-1,0,1\}$ and a prompt embedding of 00 with $N=2$. Then, the optimal cluster is either $\{-1,0\}$ or $\{0,1\}$, yet TTMM would pick $\{-1,1\}$ since its centroid averages out differences in the embeddings. Since TTMM first clusters the data, there is reason to expect that each cluster has small diameter, in which case the centroid is a good approximation of the individual data points.
@@
As a final step, to compensate for the approximation error of the chosen expert model induced by steps (1) and (2), TTMM merges multiple expert models. Intuitively, and as visualized in Figure 3, this is motivated by multiple experts being able to cover distinct features of the prompt that may not all be captured by a single expert model. We find empirically that merging a small number of experts can improve model performance without incurring a significant test-time overhead (cf. Footnote 3). TTMM computes merging coefficients via cross-attention, which is equivalent to weighting the active experts with an RBF kernel: [^4] See Appendix B for a proof.
@@ 4 Results
Llama-3.2-1B $8.674$ $2.611$
[…]
    + TTMM (1 active expert) $7.669$ $2.552$
@@
Table 1: Perplexity (lower is better) across evaluation datasets and base models with $100$ experts. Bold numbers denote the best test-time efficient method, and underlined numbers denote the best overall. TTMM outperforms standard multitask fine-tuning of a single model.
We evaluate language modeling with causal language models. We report the perplexity of the models on holdout data from the Wikipedia (English, Wikimedia Foundation, 2025) and GitHub (Python code) corpora, comprising 6.4M and 7.2M documents, respectively. As base models, we use Llama-3.2-1B (Grattafiori et al., 2024) and Qwen-2.5-1.5B (Yang et al., 2024a). For clustering and model selection, we use embeddings from the all-mpnet-base-v2 model (Reimers & Gurevych, 2019; Song et al., 2020). Following Hardt & Sun (2024) and Hübotter et al. (2025), we evaluate TTT by training a LoRA adapter for a single gradient step each on the $N=100$ nearest neighbors to the prompt embedding, most to least similar. We also compare against a single model, which is fine-tuned for a single epoch on the respective corpus. This fine-tuned model can be seen as an upper bound on the performance of approaches to multitask learning, since multitask learning typically trains one model per meta-task (i.e., per corpus) and merging models from multiple meta-tasks typically reduces performance on any individual meta-task. We will open-source our code and models for reproducing our experiments.
We show in Footnote 3 and Table 1 that TTMM with $10$ active experts achieves close to the accuracy of TTT, and much higher accuracy than a single model that was fine-tuned on the respective corpus. Instead of merging experts in parameter space, we also evaluate merging experts in prediction space by ensembling the experts. We find that this yields marginally higher accuracy, yet it has a significantly larger test-time overhead. Overall, merging experts in parameter space achieves the best trade-off between accuracy and test-time efficiency.
Figure 5: On each dataset, we evaluate the performance improvement of TTMM’s cross-attention mechanism compared to uniform weighting of active experts. Uniform weighting (dotted) selects the experts closest to the prompt embedding, and weights them equally. We see across merging (red) and ensembling (light blue) that the cross-attention mechanism outperforms uniform weighting significantly. Notably, in Wikipedia, combining multiple experts helps only when using cross-attention. The number of experts is $K=1000$.
In Figure 5, we evaluate perplexity with merging coefficients computed via cross-attention to uniform weighting of active experts. We find that cross-attention significantly outperforms uniform weighting. In particular, on the Wikipedia dataset, merging multiple experts does not help at all when weighting them uniformly. We also compare against merging coefficients, such as leveraging the logit-entropy of the experts (cf. Appendix D), but we do not find them to outperform TTMM’s cross-attention while being significantly more computationally expensive. We find that sparsity $\tau=0.01$ provides the best trade-off between accuracy and test-time efficiency, and we fix it in all other experiments. Furthermore, on our evaluation corpora, we need to select roughly 10 active experts to cover the local neighborhood of the prompt (cf. Figure 6 (right)), which is also when the performance of TTMM begins to saturate.
Refer to caption Refer to caption
Figure 6: Left: Perplexity (brighter is better) of $1000$ experts on clusters of the Wikipedia dataset. The lowest perplexity is achieved on the diagonal, i.e., when the expert model is evaluated on holdout data from the cluster it was trained on. The block-diagonal structure is an artifact of bisecting k-means, which leads to similar neighboring clusters. Middle: In our evaluation datasets, the rate of reduction in k-means loss is largest until around $100$ clusters. Right: For 100 random samples from each cluster’s holdout set, we measure whether the correct cluster is among the top-$k$ closest clusters and report the average pass@$k$ accuracy.
Insight 3: Scaling the number of experts by an order of magnitude improves accuracy.
We find, as summarized in Footnote 3, that TTMM with $100$ experts significantly outperforms TTMM with $10$ experts or the multitask fine-tuned model (i.e., a single expert). This indicates that scaling the number of experts beyond the number of meta-tasks can increase accuracy substantially. Nevertheless, we also find that scaling the number of experts to $1000$ does not further improve accuracy on our evaluation corpora. This may be because the additional experts do not capture any meaningful new information (cf. Figure 6 (middle)) or because a larger total number of experts requires more active experts which may increase model interference during merging.
Insight 4: Local experts outperform the base model in their local neighborhood.

### v2

@@ Abstract
Mixture of expert (MoE) models are a promising approach to increasing model capacity without increasing inference cost, and are core components of many state-of-the-art language models. However, current MoE models typically use only few experts due to prohibitive training and inference cost. We propose Test-Time Model Merging (TTMM) which scales the MoE paradigm to an order of magnitude more experts and uses model merging to avoid almost any test-time overhead. We show that TTMM is an approximation of test-time training (TTT), which fine-tunes an expert model for each prediction task, i.e., prompt. TTT has recently been shown to significantly improve language models, but is computationally expensive. We find that performance of TTMM improves with more experts and approaches the performance of TTT. Moreover, we find that with a 1B parameter base model, TTMM is more than 100×100$\times$100 × faster than TTT at test-time by amortizing the cost of TTT at train-time. Thus, TTMM offers a promising cost-effective approach to scale test-time training.
Figure 1: Accuracy gains of TTMM when scaling test-time compute (left) and pre-training compute (right). The normalized perplexity between base model and TTT model is averaged across evaluation datasets and models. Left: We measure language modeling ability against the time to generate 100 tokens. TTMM approaches the performance of TTT without almost any test-time overhead when increasing the number of active experts. Ensembling rather than merging experts improves performance, but at a much higher test-time cost. Standard fine-tuning on the training data improves performance without any test-time overhead, but does not yield close to the performance of TTT. Right: We measure the performance of TTMM against the total number of experts. Whereas common MoE models use few experts, TTMM scales to an order of magnitude more experts, improving performance with more experts. Scaling the number of experts in TTMM does not affect training FLOPs or required GPU memory. [^2] Scaling the number of experts increases CPU host memory, which is cheaper than GPU memory. Training the individual experts is embarrassingly parallelizable.
@@ 1 Introduction
The standard paradigm of machine learning separates training and testing. In training, we learn a model by inductively extracting general rules from data, while in testing, we apply this model to new unseen data. We study an alternative transductive paradigm called test-time training (TTT, Sun et al., 2020) where we use a specific expert model for each prediction task. Variations of this paradigm have been studied since the inception of the machine learning field (Cleveland, 1979; Cleveland & Devlin, 1988; Atkeson et al., 1997; Bottou & Vapnik, 1992). More recently, fine-tuning large pre-trained neural networks at test-time has regained interest in computer vision (e.g., Jain & Learned-Miller, 2011; Shocher et al., 2018; Sun et al., 2020; Gandelsman et al., 2021; Ruiz et al., 2023) and language modeling (Krause et al., 2018; 2019; Hardt & Sun, 2024; Sun et al., 2024; Hübotter et al., 2025). Hardt & Sun (2024) and Hübotter et al. (2025) fine-tune a pre-trained language model on data related to the prompt, and show that this can significantly improve language modeling performance. Beyond pure language modeling, TTT has been instrumental in state-of-the-art approaches to abstract reasoning (Akyürek et al., 2025), video generation (Dalal et al., 2025), and also been shown to improve reasoning language models (Zuo et al., 2025; Simonds & Yoshiyama, 2025).
@@
In our experiments, TTMM approaches the language modeling performance of TTT without incurring almost any significant compute or memory cost (cf. Footnote 2). Figure 2 illustrates the difference between TTT and TTMM.
@@
We contribute the following key findings, summarized in Footnote 2:
@@
    TTMM outperforms multitask model merging: We find that training and merging many expert models, even within a single domain such as “coding in Python”, outperforms training and merging few models specialized on broader tasks.
@@ 2 Related Work
Combining the predictions of multiple models has been a long-standing practice in machine learning, e.g., in the form of majority voting (Wang et al., 2023) or by ensembling of predictions (Dietterich, 2000; Wortsman et al., 2022), which has been used to avoid catastrophic forgetting (e.g., Bagatella et al., 2025). However, merging predictions of many neural networks is computationally inefficient since it requires a separate forward pass for each model. Instead, TTMM merges in parameter-space, and only requires a single forward pass of the merged model, resulting in a negligible increase of compute and memory cost compared to inferencing a single model. The idea of merging multiple models by averaging their parameters (not their predictions!) is grounded in the surprising phenomenon of mode-connectivity (Garipov et al., 2018; Wilson, 2025), which observes that modes in ensembles of models are connected by paths of small loss. This has motivated a variety of model merging techniques, including stochastic weight averaging (Izmailov et al., 2018) and model soups (Wortsman et al., 2022). Recently, model merging has been studied predominantly in the context of merging few expert models with the aim of retaining performance on their respective meta-tasks (Ilharco et al., 2023; Matena & Raffel, 2022; Jin et al., 2023; Ainsworth et al., 2023). Merging model parameters may lead to interference, which has been a key focus of recent work (Yadav et al., 2023; Yang et al., 2024c; Tang et al., 2024b; Daheim et al., 2024; Yu et al., 2024; Jung et al., 2024).
@@
While the above literature focuses on multitask learning with typically only a handful of expert models, TTMM clusters the training data into many local neighborhoods and trains a separate local model for each cluster. Like TTMM, Branch-Train-Merge (BTM, Li et al., 2022; Gururangan et al., 2023) clusters the training data into local neighborhoods, trains a separate local model for each cluster, and dynamically selects a subset of these models at test-time. However, unlike BTM, TTMM uses an order of magnitude more local models and merges their parameters instead of their predictions, leading to substantially faster inference. In concurrent work, Qiu et al. (2025) study TTMM in a continual learning setting.
@@
The mixture of experts (MoE, Shazeer et al., 2017) paradigm increases model capacity without increasing inference cost by introducing conditional routing of inputs to subsets of learnable parameters (i.e., “experts”). Training MoE models is challenging due to the need for joint training of the experts and the router on substantial multitask data. Moreover, MoE models evaluate all active experts individually, which can still require substantial memory. This in turn implies that the number of experts is typically limited to a small number. Recent studies have focused on challenges such as reducing synchronization (Sukhbaatar et al., 2024; Zhang et al., 2024b; 2025), load balancing of experts (Clark et al., 2022; Zhou et al., 2022), and improving expert specialization (Dai et al., 2024). In contrast, TTMM approximates the MoE paradigm by embarrassingly parallelizing the training of the local models and merging model parameters at test-time, which permits scaling to many orders of magnitude more experts and which enables approximation of the “maximally local” TTT.
@@ 3 TTMM: Test-Time Model Merging
Our proposed method, TTMM, has two main stages: train-time and test-time. At train-time, TTMM clusters the training data into many local neighborhoods and trains a separate expert model for each cluster. Then at test-time, TTMM dynamically composes these expert models to form a single task-specific model. We discuss both stages in more detail below.
@@
3.1 Latency
In contrast to MoE models, the number of experts in TTMM is typically too large for all experts to fit in GPU memory at once. For example, with Llama-3.2-1B as base model, each LoRA adapter has size approximately 172172172 MB, and with 100010001000 clusters, the expert models would require prohibitive 172172172 GB of GPU memory. Instead, we dynamically load active experts into GPU memory and merge their parameters before inference.
[…]
@@ 3.1 Latency
In Figure 4 (middle), we evaluate the practical test-time overhead of TTMM for language generation. TTMM incurs only a small fixed overhead which, when merging 101010 experts, is equal to generating around 202020 tokens. When generating more than 100100100 tokens, the overhead of TTMM compared to inference with the base model is negligible. In contrast, ensembling has a multiplicative overhead that grows with the number of active experts. Comparing to TTT, a single test-time gradient step with Llama-3.2-1B takes approximately 150150150 milliseconds (Hübotter et al., 2025). With 100100100 test-time gradient steps, the total test-time overhead (without any data selection and loading) is already around 151515 seconds. In contrast, TTMM with 101010 active experts has a constant overhead of around 115115115 milliseconds, which is a more than 125×125$\times$125 × speedup compared to TTT.
Balancing Latency and Performance.
The exact Pareto frontier of latency and performance varies depending on generation length. Footnote 2 plots the Pareto frontier for 100 generated tokens and Figure 4 (left) shows how the latency changes for other generation lengths. We do not show TTT in Figure 4 (left) since the constant overhead of TTT before generating any token is larger than the time to generate 1k tokens with TTMM (and 10 active experts).
@@ 3.2 Test-Time Model Merging approximates Test-Time Training
Suppose the neural network $f(\boldsymbol{x};\boldsymbol{\theta})$ is $L$-Lipschitz in $\boldsymbol{\theta}$ and that the loss is $G$-smooth in $\boldsymbol{x}$, with any $L,G>0$. Further, suppose we adapt models via $T\geq 1$ steps of gradient descent with step size $\eta>0$ from a shared initialization $\boldsymbol{\theta}$. For any prompt $\boldsymbol{x^{\star}}\in\mathcal{X}$ and $N\geq 1$, let $\boldsymbol{\theta}_{\boldsymbol{x^{\star}}}$ be the TTT model trained on the set $\mathcal{D}_{\boldsymbol{x^{\star}}}$ of the $N$ nearest neighbors to $\boldsymbol{\phi}(\boldsymbol{x^{\star}})$ in $\mathcal{D}$. Let $\boldsymbol{\theta}^{\prime}$ be the expert model trained on any $\mathcal{D}^{\prime}\subseteq\mathcal{D}$. Then, if $\mathcal{D}^{\prime}$ contains at least one nearest neighbor from $\mathcal{D}_{\boldsymbol{x^{\star}}}$,
$\left\|f(\boldsymbol{x^{\star}};\boldsymbol{\theta}_{\boldsymbol{x^{\star}}})-f(\boldsymbol{x^{\star}};\boldsymbol{\theta}^{\prime})\right\|\leq\eta TLG(\operatorname*{diam}(\mathcal{D}_{\boldsymbol{x^{\star}}})+\operatorname*{diam}(\mathcal{D}^{\prime})).$ (1)
To understand the intuition behind this proposition, consider an expert model trained on a cluster $\mathcal{D}^{\prime}$ that contains parts of the local neighborhood of the prompt embedding (for this, the entire clustering must cover the dataset). Then, the above proposition shows that this expert model is a good approximation of the TTT model if the diameter of $\mathcal{D}^{\prime}$ is also small. Consequently, this motivates the approximation of TTT by TTMM with many expert models covering the entire dataset, each trained on a small cluster. Training expert models on this partitioning of the dataset does not cost more FLOPs or GPU memory than training a single model on the entire dataset. This contrasts with TTT, which considers a separate expert model for each possible recombination of the training data into subsets of size $N$.
@@
[…]
@@
As a final step, to compensate for the approximation error of the chosen expert model induced by steps (1) and (2), TTMM merges multiple expert models. Intuitively, and as visualized in Figure 3, this is motivated by multiple experts being able to cover distinct features of the prompt that may not all be captured by a single expert model. We find empirically that merging a small number of experts can improve model performance without incurring a significant test-time overhead (cf. Footnote 2). TTMM computes merging coefficients via cross-attention, which is equivalent to weighting the active experts with an RBF kernel: [^3] See Appendix B for a proof.
@@ 4 Results
Llama-3.2-1B 8.6748.6748.674 2.6112.6112.611
   + fine-tuning 7.8497.8497.849 2.5812.5812.581
   + TTMM (1 active expert) 7.6697.6697.669 2.5522.5522.552
@@
   + TTT 7.231 2.177
Table 1: Perplexity (lower is better) across evaluation datasets and base models with 100100100 experts. Bold numbers denote the best test-time efficient method, and underlined numbers denote the best overall. TTMM outperforms standard multitask fine-tuning of a single model.
We evaluate language modeling with causal language models. We report the perplexity of the models on holdout data from the Wikipedia (English, Wikimedia Foundation, 2025) and GitHub (Python code) corpora, comprising 6.4M and 7.2M documents, respectively. As base models, we use Llama-3.2-1B (Grattafiori et al., 2024) and Qwen2.5-1.5B (Yang et al., 2024a). For clustering and model selection, we use embeddings from the all-mpnet-base-v2 model (Reimers & Gurevych, 2019; Song et al., 2020). Following Hardt & Sun (2024) and Hübotter et al. (2025), we evaluate TTT by training a LoRA adapter for a single gradient step each on the 100100100 nearest neighbors to the prompt embedding, most to least similar. We also compare against a single model, which is fine-tuned for a single epoch on the respective corpus. This fine-tuned model can be considered an upper bound on the performance of approaches to multitask learning, since multitask learning typically trains one model per meta-task (i.e., per corpus) and merging models from multiple meta-tasks typically reduces performance on any individual meta-task.
Our code is available at https://github.com/rbertolissi/ttmerge. We share fine-tuned baselines and TTMM-MoEs on Huggingface.
We show in Footnote 2 and Table 1 that TTMM with 101010 active experts achieves close to the accuracy of TTT, and much higher accuracy than a single model that was fine-tuned on the respective corpus. Instead of merging experts in parameter space, we also evaluate merging experts in prediction space by ensembling the experts. We find that this yields marginally higher accuracy, yet it has a significantly larger test-time overhead. Overall, merging experts in parameter space achieves the best trade-off between accuracy and test-time efficiency. Interestingly, TTMM outperforms TTT on the Wikipedia dataset. This may be because TTMM can incorporate knowledge from a larger set of experts, each trained on a few thousands of data points. In contrast, TTT is only training on the 100 most related neighbors, which may miss some information if there are more than 100 related data points. In our view, the empirical difference between TTMM and TTT is an interesting direction for further study.
Figure 5: On each dataset, we evaluate the performance improvement of TTMM’s cross-attention mechanism compared to uniform weighting of active experts. Uniform weighting (dotted) selects the experts closest to the prompt embedding, and weights them equally. We see across merging (red) and ensembling (light blue) that the cross-attention mechanism outperforms uniform weighting significantly. Notably, on Wikipedia, combining multiple experts helps only when using cross-attention. The number of experts is $K=1000$.
In Figure 5, we evaluate perplexity with merging coefficients computed via cross-attention to uniform weighting of active experts. We find that cross-attention significantly outperforms uniform weighting. In particular, on the Wikipedia dataset, merging multiple experts does not help at all when weighting them uniformly. We also compare against merging coefficients, such as leveraging the logit-entropy of the experts (cf. Section D.3), but we do not find them to outperform TTMM’s cross-attention while being significantly more computationally expensive. We find that sparsity $\tau=0.01$ provides the best trade-off between accuracy and test-time efficiency, and we fix it in all other experiments. Furthermore, on our evaluation corpora, we need to select roughly 10 active experts to cover the local neighborhood of the prompt (cf. Figure 6 (right)), which is also when the performance of TTMM begins to saturate.
Refer to caption Refer to caption
@@
Insight 3: Local experts outperform the base model in their local neighborhood.
@@
Choosing the number of (active) experts.
We find, as summarized in Footnote 2, that TTMM with 100100100 experts significantly outperforms TTMM with 101010 experts or the multitask fine-tuned model (i.e., a single expert). This indicates that scaling the number of experts beyond the number of meta-tasks can increase accuracy substantially. Nevertheless, we also find that scaling the number of experts to 100010001000 does not further improve accuracy on our evaluation corpora. In the following, we discuss considerations for the number of experts in TTMM:
  • Number of experts ($K$): The total number of experts has almost no direct effect on latency. However, more experts require a large amount of static CPU memory. Further, increasing the number of experts also increases the separation of knowledge/skills between experts. This then requires more active experts to retain the same performance, which also increases latency and may lead to model interference due to merging. To prevent knowledge fragmentation, $K$ should be determined depending on the training set. In our experiments, choosing 100 experts is indicated by the “elbow method”, common in clustering (cf. Figure 6 (middle)).
  • Number of active experts ($N$): The number of active experts has the strongest effect on latency, since all active experts have to be loaded from CPU into GPU memory before inference. Our ablation in Figure 6 (right) indicates that with 101010 active experts, the “correct” expert is likely to be active, while latency remains minimal.
Evaluation on MMLU.
We additionally evaluate TTMM on MMLU (Hendrycks et al., 2021) as a non-perplexity task. We use (non-instruct) Llama-3.2-1B as base model, which we then fine-tune for one epoch on the MMLU training set. This fine-tuned model achieves approximately the same performance as instruction-tuned Llama-3.2-1B. We use the non-instruct model to avoid any confounding effect of RL-training or instruction-tuning on privileged data from the instruct model. As embedding model, we use Stella (stella_en_400M_v5) but obtain similar results with MPNet (cf. Section D.2). We train 100 experts on the MMLU training set, using the globally fine-tuned model as initialization. We summarize the results in Table 2 and find that TTMM slightly increases accuracy. A broader evaluation on other non-perplexity tasks and non-language domains is an important direction for future work.
Model Humanities Social Sciences STEM Other Overall
Fine-tuned 44.8 54.24 40.79 54.43 48.10
TTMM (1 expert) 45.25 54.92 40.98 54.30 48.41
TTMM (3 experts) 45.33 55.02 40.98 54.39 48.48
TTMM (10 experts) 45.63 55.15 41.39 54.55 48.74
TTMM (15 experts) 45.46 55.31 41.77 55.26 48.96
Table 2: Accuracy (in %) of TTMM on MMLU. Bold numbers denote the highest accuracy.
Ablations.
We perform a series of ablations, which we summarize in Appendix D. Notably, we find that (1) larger embedding models need not lead to improved performance of TTMM and that (2) TTMM performs similarly when fixing the number of experts as to when selecting the number of active experts dynamically via thresholding.